OpenStudy (anonymous):

Explain the following in your own words: (a + b)2 ≠ a2 + b2. Give a numerical example to illustrate this.

OpenStudy (anonymous):

This is a kind of problem that you should do yourself. The words I use may not be the ones you use. Anyway, the statement says that a squared plus b squared isn't necessarily equal to the sum of a and b squared.\[3^2 + 4^2 = 25; (3+4)^2=7^2=49\]

OpenStudy (anonymous):

In words this means that the square of a sum is not equal to the sum of the individual parts squared. An example would be (3 + 2)^2 ≠ 3^2 + 2^2 as the first part is equal to 25 and the second part is equal to 13.

OpenStudy (anonymous):

Thanks!!!